SwaggyMark:

Solve each question with the quadratic formula 6a² + 10 = 0 I know the answer but I'm not to sure how imaginary roots work. I would really appreciate if someone could explain how imaginary roots work! :)

Angle:

Hello! Do you have the simplified square root before where you got stuck? / after you plug into the quadratic formula?

SwaggyMark:

\[\frac{ \sqrt{-240} }{ 12 }\]

Angle:

wonderful are you aware of how to simplify square roots without worrying about the negative?

Angle:

for example: \(\sqrt{12}=\sqrt{4*3}=2\sqrt{3}\) but you do it with \(\sqrt{240}\)

SwaggyMark:

\[\sqrt{240}=\sqrt{4*4*15}=4\sqrt{15}\]

Angle:

perfect! this means that \(\sqrt{-240}=4\sqrt{-15}\)

Angle:

now, the only thing different with the imaginary stuff is... \(\sqrt{-1}=i\) so \(\sqrt{-15}=\sqrt{-1*15}=i~\sqrt{15}\)

SwaggyMark:

would it be written as \[4i \sqrt{15}\]or\[4\sqrt{15}i\]

Angle:

for handwriting purposes, everything not under the squareroot should go in front of the squareroot, just in case you draw it too long on the paper by accident

SwaggyMark:

so you would then have \[\frac{-0± 4i \sqrt{15} }{ 12 }\] which can be simplified to \[\frac{ 0±i \sqrt{15} }{ 3 }\] which is the final answer?

Angle:

you might not need the 0 in your final answer, but that is correct :)

SwaggyMark:

thanks